Evaluate each function 3) h)4t-3; Find h(3) 4) fn)- n2-4; Find /-5)
Answer #1

3) Given, h(t) = 4t-3

Now, h(3) = 4*3-3

i.e., h(3) = 12-3

i.e., h(3) = 9

4) Given, f(n) = n2-4

Now, f(-5) = (-5)2-4

i.e., f(-5) = 25-4

i.e., f(-5) = 21